# OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation

### v1

@@ Abstract
Radiology report generation (RRG) aims to automatically produce clinically faithful reports from chest X-ray images. Prevailing work typically follows a scale-driven paradigm, by multi-stage training over large paired corpora and oversized backbones, making pipelines highly data- and compute-intensive. In this paper, we propose Oracle-educated GRPO (OraPO) with a FactScore-based reward (FactS) to tackle the RRG task under constrained budgets. OraPO enables single-stage, RL-only training by converting failed GRPO explorations on rare or difficult studies into direct preference supervision via a lightweight oracle step. FactS grounds learning in diagnostic evidence by extracting atomic clinical facts and checking entailment against ground-truth labels, yielding dense, interpretable sentence-level rewards. Together, OraPO and FactS create a compact and powerful framework that significantly improves learning efficiency on clinically challenging cases, setting the new SOTA performance on the CheXpert Plus dataset (0.341 in F1) with 2–3 orders of magnitude less training data using a small base VLM on modest hardware.
Figure 1: Illustration of mainstream data/compute-intensive pipelines (upper-left) versus our data-efficient pipeline (upper-right). Bottom: our method achieves the SOTA performance for RRG with less than 0.1% of the training samples (vs. 1.27 M) used by best-performing baselines and a much smaller model, demonstrating strong performance under tight data and compute budgets.
footnotetext: Work done during an internship at Oracle Health & AI. Joint senior authors.
@@ 1 Introduction
Recently, Group Relative Policy Optimisation (GRPO) has emerged as a strong and compute-efficient RL-based learning algorithm for many hard problems with verifiable rewards. Its efficiency comes from two fronts: (i) lower memory/compute by dispensing with the trainable value critic and instead computing group-normalised advantages from the relative rewards of multiple samples [45, 31]; and (ii) a simplified training recipe that enables “RL-only” regimes (i.e., without costly SFT/alignment), avoiding massive pre-training/alignment corpora [54, 16, 10].
@@
However, applying GRPO to the RRG task reveals two challenges. Challenge (i): we observe that vanilla GRPO often produces all-zero reward groups early in training (e.g. 30% in the first 50 steps as shown in Fig. 2 left), causing vanishing gradients and wasted rollouts. Recent variants attempt to fix this by resampling until a non-zero reward appears [75] or by enlarging the group size [55], both of which are still suboptimal due to increased cost of more rollouts. Another line interleaves SFT with RL to maintain gradient flow, yet still leaves low-quality rollouts unused, wasting costly sampling [40, 11, 76, 39, 37]. Challenge (ii): vanilla GRPO requires readily verifiable scalar rewards. Unlike multiple-choice QA, maths, or coding, where the reward is typically a binary final-answer check, radiology reports are long-form, multi-fact, and must be consistent across sentences, making reward design both critical and difficult. In practice, RLHF-style RRG work often uses reference-overlap metrics that mainly capture fluency/surface similarity (e.g. BLEU/CIDEr [33, 47]) or a report-level clinical metric [80, 44, 74]. These proxies under-penalise sentence-level factual errors and cross-sentence contradictions, yielding fluent yet clinically misleading narratives (missed positives, unsupported claims). Hence, a reward must be able to assess factual correctness of each statement of a radiology report in an efficient manner.
@@
In this paper, we propose Oracle-educated Group Relative Policy Optimisation: OraPO, a novel RL-based learning algorithm that tackles RRG’s challenges in a data- and compute-efficient manner. OraPO addresses Challenge (i) of GRPO by dynamically injecting oracle supervision (i.e. ground-truth) for groups with all-zero rewards and teaching the policy to avoid repeating low-quality generations. Specifically, When a sampled group yields all-zero rewards (group 1 in Fig. OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation, upper-right), OraPO triggers a lightweight Direct Preference Optimisation (DPO) update that directly prefers the ground-truth report over the zero-reward rollouts as ready-made negatives, without extra compute or annotations. This converts failed exploration on difficult cases (hard or low-prevalence studies) into useful gradients, moving the policy away from low-quality modes and toward the ground-truth. We also introduce a simple adaptive mixing weight to increase DPO’s influence when zero-reward groups are frequent, which gradually tapers once GRPO provides informative advantages, making OraPO salvage otherwise wasted batches, stabilise training, and accelerate convergence.
@@
Moreover, to provide OraPO with clinically faithful supervision, we introduce a FactScore-based reward [42] designed to address Challenge (ii). We treat the generated report as its own rationale, extract atomic clinical facts, and test their entailment against the ground-truth label set. Facts that support a label contribute positively; unsupported or contradictory labels incur penalties, yielding a dense, interpretable per-label reward. This stands in sharp contrast to the original GRPO reward, which assigns reward solely to the final answer while treating the reasoning trace largely as a byproduct of generation, without ensuring it is factually grounded or consistently aligned with the answer. Our reward design ensures the generated report (serving in lieu of reasoning trace) not only produces correct cognitive reasoning trace similar to that of an experienced radiologist but also yields correct diagnostic findings, which is crucial in domains such as healthcare in which safety is of paramount importance.
@@
    We present OraPO, an Oracle (DPO) educated GRPO that turns failed explorations into direct preference supervision and uses an adaptive mixing schedule to emphasize DPO when GRPO’s signal vanishes, significantly improving data- and compute-efficiency. To the best of our knowledge, this is the first work that integrates direct preference learning with GRPO-based RL for efficient training.
@@
    OraPO sets the new SOTA performance for RRG, surpassing the previous best model in terms of clinical effectiveness F1 score. This is achieved while being exceptionally data-efficient, with 2–3 orders of magnitude less training data. OraPO is trained on only 1K samples. In contrast, the current best model MambaXray-L is trained on 1.27M samples and all the other models are trained on at least 223K samples. At the same time, OraPO is also compute-efficient: it is based on QWen2.5-VL-3B, a moderately-sized VLM on modest hardware (4x A10 GPUs).
@@ 2 Related Work
Automatic Radiology Report Generation (RRG) aims to turn imaging studies (e.g., chest X-rays) into clinically usable free-text reports. Early approaches adopted seq2seq networks (e.g., CNN–RNN and Transformer) [28, 13] for training on large paired corpora such as MIMIC-CXR [29]. Recent work advances three fronts: (i) knowledge- and rule-guided generation, which injects medical priors to curb hallucination [6] and performs disease-aware image–text alignment for trustworthy text [50]; (ii) leveraging prior studies and multi-view images to improve coverage and consistency [22, 36]; and (iii) scaling/standardizing with pre-training, releasing multi-stage pre-training recipes [62, 22]. In parallel, LLM-driven RRG explores instruction-tuned or preference-aligned training [35, 32]. Despite progress, these approaches often rely on large paired corpora and/or costly computation resources, leading to deficiency.
Group Relative Policy Optimisation (GRPO) is a critic-free, reinforcement learning objective for generative LLMs, introduced in DeepSeekMath [54]. It samples multiple completions per prompt, computes group-normalised advantages from relative rewards, and updates the policy with clipped ratios to a frozen reference; removing the critic reduces memory/compute versus PPO while preserving stability [54, 21, 45, 31]. Recent variants mainly target optimisation stability: DR.GRPO [38] mitigates length bias and skewed group rewards; CISPO [9] clips importance-sampling weights; GSPO [77] adopts sequence-level clipping for MoE; DAPO [75] resamples until a non-zero reward appears. However, these methods were not designed to improve data efficiency: most target optimisation stability while leaving data volume and training epochs essentially unchanged, and some of them often increase sampling/compute [75, 59, 55].
Direct Preference Optimisation (DPO) trains a policy from preference pairs to prefer the chosen response over the rejected one [79]. Recent variants streamline or robustify this recipe: SimPO [41] applies length-normalised mean log-probabilities with a simplified reference; Cal-DPO [67] calibrates the implicit reward to human scales; ORPO [24] drops the reference via an odds-ratio objective integrated into SFT; and KTO [29] learns from unary thumbs-up/down signals instead of pairs. In our RRG framework, DPO naturally serves as an oracle step inside OraPO to convert those failed GRPO rollouts into useful preference updates against the ground-truth, leading to data- and compute-efficient learning.
@@ 3 Methodology
In this section, we first introduce the background of Group Relative Policy Optimisation (GRPO) and Direct Preference Optimisation (DPO) (§3.1). We then identify a failure mode of GRPO in radiology report generation (RRG), when a pretrained model produces many low-quality groups (e.g., zero rewards) that yield little or no gradient. This observation motivates our proposd OraPO algorithm, which efficiently converts such failed explorations into preference supervision (§3.2). Finally, we design the FactS Reward, a fact-level entailment signal that aligns reports with diagnostic labels and provides dense, interpretable feedback for reinforcement learning (§3.3).
@@ 3.2 OraPO: Oracle-educated GRPO
Exploration Failure of GRPO in RRG. On radiology report generation, applying vanilla GRPO to a base VLM with limited radiology knowledge often fails to explore productively. The root reason is that the base model carries very limited domain knowledge about radiology. Most base VLMs are pre-trained on general vision–language corpora and tasks, for example, everyday-image captioning on MS COCO [34] and open-domain visual question answering on VQA v2 [19], rather than medical imaging and reporting. Hence, it tends to miss or misstate rare and/or clinically challenging findings (e.g., small pneumothorax, subtle interstitial edema, or linear atelectasis) when asked to produce clinical narratives. For many chest X-ray inputs, the sampled group of completions is uniformly low quality with near-zero rewards. Thus, the policy receives little learning signal to drive effective exploration.
@@
Intuitively, a high ZRR flags failure cases where preference education must lead. OraPO responds by increasing $w_{i}^{(t)}$ and applying DPO updates that teach the policy to reject the sampled negatives and align with the ground-truth report. This converts groups into labeled preferences and supplies gradients exactly when GRPO has none/few. A low ZRR indicates that GRPO is uncovering usable reward. This can happen on simpler studies, and it also emerges after DPO has taught the model the cues for difficult findings (e.g., small apical pneumothorax or subtle interstitial edema), enabling a shift from education to exploration and exploitation. In practice, we observe the intended training dynamic, as shown in Fig. 2: early training is education-heavy with larger ratio of zero-reward groups $\rightarrow w^{(t)}$, then $w^{(t)}$ gradually declines as the policy learns and GRPO takes over.
@@ 3.3 FactScore-based Reward Design
Table 1: Experimental results on the CheXpert Plus dataset [8] using RRG task-specific algorithms. We report per-label Precision, Recall, and F1 (14 pathologies). Train Size is the number of training samples. Some baselines train on multiple corpora and/or in multiple stages. Best and second best are in bold and underline, respectively.
@@
Table 2: Experimental Results on the CheXpert Plus benchmark [8] across diverse LLMs/VLMs, all supervised fine-tuned with R2GenGPT [65] on CheXpert Plus. We report macro Precision, Recall, and F1 (14 pathologies). The Params listed denotes the parameters that need to be tuned in the training phase. FT-Size indicates the number of samples used for fine-tuning. Best and second best are in bold and underline, respectively.
@@
Table 3: Ablation on CheXpert Plus [8]: Impact of the FactS reward and OraPO. The first row indicates the base Qwen2.5-VL-3B-Instruct [4] direct test results, without any fine-tuning. The second row corresponds to a naïve GRPO-based RLHF baseline using accuracy as the reward [21]. We report per-label Precision, Recall, and F1 across 14 findings.
FactS OraPO Train Size Precision, Recall, F1
$\times$ $\times$ 0 0.097,  0.104,  0.034
$\times$ $\times$ 1, 000 0.026,  0.162,  0.089
$\checkmark$ $\times$ 1, 000 0.204,  0.605,  0.291
$\checkmark$ $\checkmark$ 400 0.217,  0.732,  0.296
$\checkmark$ $\checkmark$ 1, 000 0.237,  0.832,  0.341

### v2

@@ Abstract
Radiology report generation (RRG) aims to automatically produce clinically faithful reports from chest X-ray images. Prevailing work typically follows a scale-driven paradigm, by multi-stage training over large paired corpora and oversized backbones, making pipelines highly data- and compute-intensive. In this paper, we propose Oracle-educated GRPO (OraPO) with a FactScore-based reward (FactS) to tackle the RRG task under constrained budgets. OraPO enables single-stage, RL-only training by converting failed GRPO explorations on rare or difficult studies into direct preference supervision via a lightweight oracle step. FactS grounds learning in diagnostic evidence by extracting atomic clinical facts and checking entailment against ground-truth labels, yielding dense, interpretable sentence-level rewards. Together, OraPO and FactS create a compact and powerful framework that significantly improves learning efficiency on clinically challenging cases, setting the new SOTA performance on the CheXpert Plus dataset (0.341 in F1) and MIMIC-CXR dataset (0.357 in F1) with 2–3 orders of magnitude less training data using a small base vision-language model on modest hardware.
Figure 1: Illustration of mainstream data/compute-intensive pipelines (upper-left) versus our data-efficient pipeline (upper-right). Bottom: on CheXpert Plus [8], our method achieves the SOTA performance for RRG with less than 0.1% of the training samples (vs. 1.27 M) used by best-performing baselines and a much smaller model, demonstrating strong performance under tight data and compute budgets.
@@ 1 Introduction
Recently, Group Relative Policy Optimisation (GRPO) has emerged as a strong and compute-efficient RL-based learning algorithm for many hard problems with verifiable rewards. Its efficiency stems from two fronts: (i) lower memory/compute by dispensing with the trainable value critic and instead computing group-normalised advantages from the relative rewards of multiple samples [63, 44]; and (ii) a simplified training recipe that enables “RL-only” regimes (i.e., without costly SFT/alignment), avoiding massive pre-training/alignment corpora [79, 114, 20, 10].
@@
However, applying GRPO to the RRG task reveals two challenges. Challenge (i): we observe that vanilla GRPO often produces all-zero reward groups early in training (e.g. 30% in the first 50 steps, as shown in Fig. 2 left), causing vanishing gradients and wasted rollouts. Recent variants attempt to fix this by resampling until a non-zero reward appears [110] or by enlarging the group size [81], both of which are still suboptimal due to increased cost of more rollouts. Another line interleaves SFT with RL to maintain gradient flow, yet still leaves low-quality rollouts unused, wasting costly sampling [58, 11, 112, 57, 54]. Challenge (ii): vanilla GRPO requires readily verifiable scalar rewards. Unlike in multiple-choice QA, maths, or coding, where the reward is typically a binary final-answer check, radiology reports are long-form, multi-fact, and must be consistent across sentences, making reward design both critical and difficult. In practice, RLHF-style RRG work often uses reference-overlap metrics that mainly capture fluency/surface similarity (e.g. BLEU/CIDEr [48, 67]) or a report-level clinical metric [120, 62, 107]. These proxies under-penalise sentence-level factual errors and cross-sentence contradictions, yielding fluent yet clinically misleading narratives (missed positives, unsupported claims). Hence, a reward must be able to assess factual correctness of each statement of a radiology report in an efficient manner.
@@
In this paper, we propose Oracle-educated Group Relative Policy Optimisation: OraPO, a novel RL-based learning algorithm that tackles RRG’s challenges in a data- and compute-efficient manner. OraPO addresses Challenge (i) of GRPO by dynamically injecting oracle supervision (i.e. ground-truth) for groups with all-zero rewards and teaching the policy to avoid repeating low-quality generations. Specifically, When a sampled group yields all-zero rewards (group 1 in Fig. OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation, upper-right), OraPO triggers a lightweight Direct Preference Optimisation (DPO) update that directly prefers the ground-truth report over the zero-reward rollouts, effectively reusing those failed rollouts as ready-made negative examples, with no extra compute or annotation needed. This converts failed exploration in difficult cases (hard or low-prevalence studies) into useful gradients, moving the policy away from low-quality modes and toward the ground-truth. We also introduce a simple adaptive mixing weight to increase DPO’s influence when zero-reward groups are frequent, which gradually tapers once GRPO provides informative advantages, thereby allowing OraPO to salvage otherwise wasted batches, stabilise training, and accelerate convergence. .
@@
Moreover, to provide OraPO with clinically faithful supervision, we introduce a FactScore-based reward [60] designed to address Challenge (ii). We treat the generated report as its own rationale, extract atomic clinical facts, and test their entailment against the ground-truth label set. Facts that support a label contribute positively; unsupported or contradictory labels incur penalties, yielding a dense, interpretable per-label reward. This stands in sharp contrast to the original GRPO reward, which assigns reward solely to the final answer while treating the reasoning trace largely as a byproduct of generation, without ensuring it is factually grounded or consistently aligned with the answer. Our reward design ensures the generated report (serving in lieu of reasoning trace) not only produces a correct cognitive reasoning trace similar to that of an experienced radiologist but also yields correct diagnostic findings, which is crucial in domains such as healthcare in which safety is of paramount importance.
Extensive experiments are conducted on CheXpert Plus [8] and MIMIC-CXR [8] datasets, where OraPO sets the new SOTA performance for RRG, surpassing the previous best model with F1 score and notably achieving 160.8% improvement in recall. This higher recall is clinically meaningful. In automated RRG tasks, recall is often prioritised over precision, as false negatives (missed abnormalities) can delay or prevent timely treatment, whereas false positives generally result only in additional clinical review [62, 78, 76, 69]. Previous studies [43, 107] emphasise maintaining high recall to avoid critical finding omissions, recognising that such oversights carry more serious clinical consequences than reduced precision. The proposed method achieves this clinical effectiveness while being exceptionally data-efficient, using 2–3 orders of magnitude less training data (i.e. trained on only 1K samples compared to the current best baseline MambaXray-L’s 1.27M samples and other baselines’ minimum of 223K samples). At the same time, OraPO is compute-efficient: it is based on Qwen2.5-VL-3B, a small-sized VLM on modest hardware (4× A10 GPUs).
@@
    We present OraPO, an Oracle (DPO) educated GRPO that turns failed explorations into direct preference supervision significantly improving data- and compute-efficiency. To the best of our knowledge, this is the first work to integrate direct preference learning with GRPO-based RL for efficient training.
@@
    We show that OraPO achieves new SOTA performance for RRG, surpassing the previous best model (MambaXray-L) in clinical effectiveness F1 and recall while using 2–3 orders of magnitude less training data.
@@ 2 Related Work
Automatic Radiology Report Generation (RRG) aims to turn imaging studies (e.g., chest X-rays) into clinically usable free-text reports. Early approaches adopted seq2seq networks (e.g., CNN–RNN and Transformer) [40, 14, 49, 109, 108] for training on large paired corpora such as MIMIC-CXR [41]. Recent work advances three fronts: (i) knowledge- and rule-guided generation, which injects medical priors to curb hallucination [6] and performs disease-aware image–text alignment for trustworthy text [72, 38]; (ii) leveraging prior studies and multi-view images to improve coverage and consistency [30, 53, 17]; and (iii) scaling/standardizing with pre-training, releasing multi-stage pre-training recipes [92, 30, 16, 18, 37]. In parallel, LLM-driven RRG explores instruction-tuned or preference-aligned training [52, 85, 45]. Despite progress, these approaches often rely on large paired corpora and/or costly computation resources, leading to deficiencies in data efficiency and practical scalability.
Group Relative Policy Optimisation (GRPO) is a critic-free, reinforcement learning objective for generative LLMs, introduced in DeepSeekMath [79]. It samples multiple completions per prompt, computes group-normalised advantages from relative rewards, and updates the policy with clipped ratios to a frozen reference; removing the critic reduces memory/compute versus PPO while preserving stability [79, 29, 113, 63, 44]. Recent variants mainly target optimisation stability: DR.GRPO [56] mitigates length bias and skewed group rewards; CISPO [9] clips importance-sampling weights; GSPO [115] adopts sequence-level clipping for MoE; DAPO [110] resamples until a non-zero reward appears. However, these methods were not designed to improve data efficiency: most target optimisation stability while leaving data volume and training epochs essentially unchanged, and some of them often increase sampling/compute [110, 89, 19, 81].
Direct Preference Optimisation (DPO) trains a policy from preference pairs to prefer the chosen response over the rejected one [119]. Recent variants [111] streamline or improve the robustness of this recipe: SimPO [59] applies length-normalised mean log-probabilities with a simplified reference; Cal-DPO [98] calibrates the implicit reward to human scales; ORPO [33] drops the reference via an odds-ratio objective integrated into SFT; and KTO [25] learns from unary thumbs-up/down signals instead of pairs. In our RRG framework, DPO naturally serves as an oracle step inside OraPO to convert those failed GRPO rollouts into useful preference updates against the ground-truth, leading to data- and compute-efficient learning.
@@ 3 Methodology
In this section, we first introduce the background of Group Relative Policy Optimisation (GRPO) and Direct Preference Optimisation (DPO) (§3.1). We then identify a failure mode of GRPO in radiology report generation (RRG), when a pretrained model produces many low-quality groups (e.g., zero rewards) that yield little or no gradient. This observation motivates our proposed OraPO algorithm, which efficiently converts such failed explorations into preference supervision (§3.2). Finally, we design the FactS Reward, a fact-level entailment signal that aligns reports with diagnostic labels and provides dense, interpretable feedback for reinforcement learning (§3.3).
@@ 3.2 OraPO: Oracle-educated GRPO
Exploration Failure of GRPO in RRG. On radiology report generation, applying vanilla GRPO to a base VLM with limited radiology knowledge often fails to explore effectively. The root reason is that the base model carries very limited domain knowledge about radiology. Most base VLMs [117] are pre-trained on general vision–language corpora and tasks, for example, everyday-image captioning on MS COCO [50, 96] and open-domain visual question answering on VQA v2 [26], rather than medical imaging and reporting. Hence, it tends to miss or misstate rare and clinically challenging findings (e.g., small pneumothorax, subtle interstitial edema, or linear atelectasis) when asked to produce clinical narratives. For many chest X-ray inputs, the sampled group of completions is uniformly low quality with near-zero rewards. Thus, the policy receives little learning signal to drive effective exploration.
@@
Intuitively, a high ZRR flags failure cases where preference education must lead. OraPO responds by increasing $w_{i}^{(t)}$ and applying DPO updates that teach the policy to reject the sampled negatives and align with the ground-truth report. This converts groups into labeled preferences and supplies gradients exactly when GRPO has none/few. A low ZRR indicates that GRPO is uncovering usable reward. This can happen on simpler studies, and it also emerges after DPO has taught the model the cues for difficult findings (e.g., small apical pneumothorax or subtle interstitial edema), enabling a shift from education to exploration and exploitation. In practice, we observe the intended training dynamic, as shown in Fig. 2: early training is education-heavy with larger ratio of zero-reward groups, higher $w^{(t)}$ is implied, then $w^{(t)}$ gradually declines as the policy learns and GRPO takes over.
@@ 3.3 FactScore-based Reward Design
Table 1: Experimental results on the CheXpert Plus dataset [8] using task-specific RRG algorithms. We report macro-averaged Precision, Recall, and F1 across 14 CheXpert pathologies. Train Size is the number of training samples. Some baselines train on multiple corpora and/or in multiple stages. Best and second best are in bold and underline, respectively.
@@
Table 2: Experimental results on the CheXpert Plus benchmark [8] across diverse LLMs/VLMs, all supervised fine-tuned with R2GenGPT [94] on CheXpert Plus. We report macro-averaged Precision, Recall, and F1 across 14 CheXpert pathologies. The Params listed denotes the parameters that need to be tuned in the training phase. FT-Size indicates the number of samples used for fine-tuning. Best and second best are in bold and underline, respectively.
@@
Table 3: Experimental results on the MIMIC-CXR dataset [8]. We report macro-averaged Precision, Recall, and F1 across 14 CheXpert pathologies. Train Size is the number of training samples. Some baselines train on multiple corpora and/or in multiple stages. Best are in bold.
Algorithm Venue Precision, Recall, F1 Train Size
R2Gen [14] EMNLP20 0.333, 0.273, 0.276 227K
CXR-RePaiR [24] ML4H21 – , – , 0.274 227K
MRG/SILC [51] TIP24 0.457, 0.337, 0.330 227K
CXRMate [65] IMU24 0.277, 0.351, 0.283 227K
HERGen [86] ECCV24 0.415, 0.301, 0.317 145K
ChEX [64] ECCV24 – , – , 0.326 227K
MambaXray-L [92] CVPR25 0.371, 0.321, 0.340 1.27M
GIT-CXR [82] INFO25 – , – , 0.327 227K
MCA-RG [100] MICCAI25 0.443, 0.306, 0.335 227K
Ours 0.242, 0.891, 0.357 1K
Table 4: Ablation on CheXpert Plus [8]: Impact of the FactS reward and OraPO. The first row indicates the base Qwen2.5-VL-3B-Instruct [4] direct test results, without any fine-tuning. The second row corresponds to a naïve GRPO-based RLHF baseline using accuracy as the reward [29]. We report per-label Precision, Recall, and F1 across 14 findings. Best are in bold.
FactS GRPO DPO SFT Train Size Precision, Recall, F1
$\times$ $\times$ $\times$ $\times$ 0 0.097,  0.104,  0.034
$\times$ $\checkmark$ $\times$ $\times$ 1,000 0.026,  0.162,  0.089
$\checkmark$ $\checkmark$ $\times$ $\times$ 1,000 0.204,  0.605,  0.291
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 400 0.217,  0.732,  0.296
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ 1,000 0.171,  0.176,  0.106
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 1,000 0.237,  0.832,  0.341
Table 5: Experimental results on CheXpert validation set with gold labels from certified radiologists [36].
Method Train Size Precision, Recall, F1
GPT4.1 [70] 0.292, 0.305, 0.253
GPT5 (Thinking) [71] 0.261, 0.421, 0.301
CheXagent [15] 8.5M 0.462, 0.226, 0.254
MambaXray-B [92] 1.27M 0.362, 0.254, 0.238
MambaXray-L [92] 1.27M 0.395, 0.326, 0.280
Ours 1K 0.234, 0.641, 0.288